# AgentHub: A Research Agenda for Agent Sharing Infrastructure

### v1

@@ Abstract.
LLM-based agents are rapidly proliferating, yet the infrastructure for discovering, evaluating, and governing them remains fragmented compared to mature ecosystems like software package registries (e.g., npm) and model hubs (e.g., Hugging Face). Recent research and engineering works have begun to consider the requisite infrastructure, but so far they focus narrowly—on distribution, naming, or protocol negotiation. However, considering broader software engineering requirements would improve open-source distribution and ease reuse. We therefore propose AgentHub, a research agenda for agent sharing. By framing the key challenges of capability clarity, lifecycle transparency, interoperability, governance, security, and workflow integration, AgentHub charts a community-wide agenda for building reliable and scalable agent ecosystems. Our vision is a future where agents can be shared, trusted, and composed as seamlessly as today’s software libraries.
@@ 1. Introduction
LLM-based agents are rapidly entering workflows, from scientific discovery (gottweis_towards_2025, ) to software engineering (zhang2024autocoderover, ). Unlike static software packages or pretrained models (JiangPTMReuse, ), Agents act with autonomy, compose tools dynamically, evolve (self-refine) over time, and can operate at scale (GoogleCloudAIAgents, ; He2025LLMMultiAgent, ). We believe that these attributes necessitate a new approach to sharing and composing the associated artifacts. As agent adoption grows, the lack of suitable infrastructure risks limiting both research progress and real-world impact.
In designing a registry for agent sharing, we can learn from the ecosystems for earlier kinds of software. Conventional software registries such as PyPI, npm, and Maven Central show the value of structured metadata, dependency graphs, and signed provenance (Zimmermann2019, ; schorlemmer2024signing, ). More recently, registries for pre-trained models, e.g., Hugging Face, expose artifacts and informal model cards, but in an effort to keep up with rapid change they omit normalized dependency and capability schemas, hampering reuse (JiangPTMReuse, ; jiang2024peatmoss, ; Jiang2025Naming, ). Meanwhile, emerging agent protocols, including the Model Context Protocol (MCP) (mcp-registry-repo, ; mcp-registry-blog, ), and the Agent Name Service (ANS) (ans, ), strengthen connectivity and naming but stop short of delivering a registry layer. The result: a fragmented landscape lacking features such as capabilities evidence and lifecycle status.
@@
We envision AgentHub, a registry to support the production and consumption of software agents (Figure 1). An agent registry shares some requirements of earlier registries. However, unlike conventional artifacts, agents’ autonomy and dynamic composition demand transparent capability schemas and evidence (section 3.2.1); their continual evolution requires lifecycle visibility and fast revocation (section 3.2.2); cross-protocol operation calls for ecosystem interoperability (section 3.2.3); and large-scale, automated reuse raises openness, governance, and security concerns (sections 3.2.4 and 3.2.5).
@@ 2.1. Lessons from Pre-Agent Software Registries
Metadata: Programming language package registries such as npm, PyPI, Maven Central, and CRAN demonstrate the importance of structured metadata and explicit dependency graphs. Package manifests (e.g., package.json, POM.xml) encode versioning, licensing, and dependency constraints in machine-readable form, enabling automated resolution and reproducible builds (pypi-pep458, ; npm-trusted, ; hf-modelcards, ). SBOM standards require explicit declaration of components and dependency relationships for provenance and traceability (ntia-sbom-minimum, ; ntia-sbom-framing, ); emerging AI/ML BOMs extend the same idea to models, datasets, and configurations (xia2024trustsoftwaresupplychains, ; bennet2025SPDXAIBOM, ). Hugging Face relies primarily on model cards with limited dependency information (JiangPTMReuse, ). For example, while some cards reference required libraries, pretrained checkpoints, or paired datasets, these links are neither mandatory nor normalized into a dependency graph schema. This lack of standardized schemas leads to inconsistent naming and hampers automated reproducibility (Jiang2025Naming, ; JiangScored22, ; jiang2024peatmoss, ). The lesson for AgentHub is that agent metadata must go beyond identifiers to include standardized schemas that capture capabilities, input–output modalities, protocol requirements, and declared dependencies.
@@
Provenance: Trust and provenance mechanisms are central to registry design. Maven Central requires every artifact to be signed with a PGP key; npm supports registry signatures and “trusted publishing” using OIDC (npm-trusted, ; npm-signatures, ); and PyPI is adopting The Update Framework (TUF) for signed metadata (pypi-pep458, ). For agents, provenance is especially critical because dynamic tool bindings and environment access amplify the risks of impersonation or poisoning. Accordingly, AgentHub should require signed metadata and reproducibility attestations for all entries.
@@
Governance and Submission Policies: Governance models highlight trade-offs between openness and safety. Npm and PyPI accept broad participation with light pre-checks, while CRAN and app stores such as Apple’s App Store impose strict manual review  (apple-appstore-guidelines, ; npm-2fa, ). Ecosystems also implement revocation: app stores can remotely disable malicious apps, and PyPI can yank faulty releases (pypi-pep458, ). Notably, the leftpad incident in npm illustrated the ecosystem-wide disruption that can follow from a poorly governed removal (npm-leftpad, ; theregister-leftpad, ). Governance requires care and is non-obvious with autonomous agents.
@@
Discovery and Quality Signals: Registries, model hubs, and extension marketplaces provide user ratings, download counts, verification badges, and metadata-rich model cards (hf-modelcards, ; hf-popularity, ). These signals allow users to identify reputable contributions at scale. AgentHub should adopt reputation systems such as usage statistics, audits, or benchmark results to complement technical metadata.
@@ 3. Research Agenda
We distinguish challenges shared by all registries (section 3.1) from those unique to agents (section 3.2).
@@ 3.2. Agent-Specific Considerations
3.2.1. Capability Clarity and Evidence
Fragmentation and Duplication: As npm’s “micro-packages” created brittle dependency chains (kula2017impactmicropackagesempiricalstudy, ; Abdalkareem2017TrivialPackages, ), agent ecosystems may spawn “micro-agents” with overlapping functions. Because agents can autonomously query registries and compose others, duplication and fragility can spread dynamically, enlarging the attack surface and degrading reliability, especially at scale as agents evolve and reimplement overlapping capabilities.
Equivalence: Beyond functional duplication, a deeper challenge is determining when two agents are truly equivalent. At a syntactic level, the same agent may appear across multiple repositories, creating confusion about which copy is authoritative. At a semantic level, multiple agents may claim the same capabilities but diverge in behavior due to nondeterminism and evolution in models, changes in context, and even hardware selection. Addressing this requires more than metadata alignment: AgentHub should support persistent identifiers, cross-registry attestations, and re-executable evidence pipelines to assess whether two agents are truly equivalent.
Abuse of Ambiguity: Ambiguity and missing metadata enable attacks in software registries (JiangPTMReuse, ); agents face the same risk. Unclear manifests let adversaries mimic popular entries or collaborators, echoing typosquatting and account hijacks in npm (Zimmermann2019, ). Agents’ autonomous installation/generation can accelerate such propagation unless strong signing and evidence requirements are enforced.
Near-term priorities are (i) standardizing machine-readable capability schemas covering capabilities, modalities, protocol roles, and dependencies enforced at publish time (JiangPTMReuse, ; Jiang2025Naming, ); (ii) supporting lightweight, re-executable (idempotent) evidence pipelines linking claims to traces or benchmark runs across versions; and (iii) adding badges or similar checks as digital nudges (brown2019nudges, ).
Some further discussion from a security view is in section 3.2.5.
@@ 3.2.2. Lifecycle Transparency
Dependency Concentration and Abandonment: The leftpad incident showed how removing even trivial packages disrupted thousands of projects. Zimmermann et al. found that a small number of npm maintainer accounts control most packages (Zimmermann2019, ), while Zerouali et al. note that popularity often masks inactivity (Zerouali2019, ). Agents face parallel risks: if many depend on a single base agent or are generated by one entity, failure or compromise could cascade broadly. Abandonment may occur not only when humans leave but also when autonomous agents stop updating, leaving stale yet discoverable entries in circulation.

### v2

# AgentHub: A Registry for Discoverable, Verifiable, and Reproducible AI Agents

@@ Abstract.
LLM-based agents are rapidly proliferating, yet the infrastructure for discovering, evaluating, and governing them remains fragmented compared to mature ecosystems like software package registries (e.g., npm) and model hubs (e.g., Hugging Face). Existing efforts typically address naming, distribution, or protocol descriptors, but stop short of providing a registry layer that makes agents discoverable, comparable, and governable under automated reuse.
We present AgentHub, a registry layer and accompanying research agenda for agent sharing that targets discovery and workflow integration, trust and security, openness and governance, ecosystem interoperability, lifecycle transparency, and capability clarity with evidence. We describe a reference prototype that implements a canonical manifest with publish-time validation, version-bound evidence records linked to auditable artifacts, and an append-only lifecycle event log whose states are respected by default in search and resolution. We also provide initial discovery results using an LLM-as-judge recommendation pipeline, showing how structured contracts and evidence improve intent-accurate retrieval beyond keyword-driven discovery. AgentHub aims to provide a common substrate for building reliable, reusable agent ecosystems.
@@ 1. Introduction
LLM-based agents are rapidly entering workflows, from scientific discovery (gottweis_towards_2025, ) to software engineering (zhang2024autocoderover, ). Unlike static software packages or pretrained models (JiangPTMReuse, ), agents act with autonomy, compose tools dynamically, evolve (self-refine) over time, and can operate at scale (GoogleCloudAIAgents, ; He2025LLMMultiAgent, ). We believe these attributes necessitate a new approach to sharing and composing the associated artifacts. As agent adoption grows, the lack of suitable infrastructure risks limiting both research progress and real-world impact.
In designing a registry for agent sharing, we can learn from the ecosystems for earlier kinds of software. Conventional registries such as PyPI, npm, and Maven Central show the value of structured metadata, dependency graphs, and signed provenance (Zimmermann2019, ; schorlemmer2024signing, ). More recently, registries for pre-trained models, e.g., Hugging Face, expose artifacts and informal model cards, but in an effort to keep up with rapid change they omit normalized dependency and capability schemas, hampering reuse (JiangPTMReuse, ; jiang2024peatmoss, ; Jiang2025Naming, ). Meanwhile, emerging agent protocols, including the Model Context Protocol (MCP) (mcp-registry-repo, ; mcp-registry-blog, ) and the Agent Name Service (ANS) (ans, ), strengthen connectivity and naming but stop short of delivering a registry layer. The result is a fragmented landscape lacking features such as capability evidence and lifecycle status.
@@
We envision AgentHub as a registry layer that sits between emerging agent protocols (e.g., MCP, A2A, ANS) and the workflows that publish, discover, and reuse agents (Figure 1). AgentHub shares baseline registry concerns with earlier ecosystems, such as namespaces, versioning, integrity, and governance. However, agents’ autonomy, dynamic composition, continual evolution, and scale impose requirements that are difficult to treat as documentation problems. Automatable reuse requires machine-checkable capability schemas and evidence (section 3.2.6), and continual evolution requires lifecycle visibility and fast revocation (section 3.2.5). Cross-protocol composition requires interoperability across descriptor dialects (section 3.2.4). Finally, scaling publishers and consumers beyond humans raises governance and security requirements that must be enforceable by default (sections 3.2.3 and 3.2.2).
This paper contributes an agenda and a substrate for testing it. We (i) distill lessons from mature software registries and emerging agent directory/protocol efforts into six requirements for agent-sharing infrastructure (Figure 2); (ii) describe an open-source reference implementation that operationalizes these requirements via a canonical manifest, version-bound evidence records, and an append-only lifecycle log, with explicit interfaces for search, resolution, and verification (section 4); and (iii) report initial discovery results and outline a broader evaluation program that targets each requirement, including deployability costs such as latency, storage growth, and scalability (sections 5 to 6). Our aim is to make the design space for agent sharing infrastructure explicit, testable, and iteratively refinable by the community; we submit this work-in-progress to JAWS to solicit feedback on whether the six requirements and the prototype’s manifest/evidence/lifecycle contract capture the right minimal core for safe, automatable agent reuse.
@@ 2.1.1. Metadata
Programming language package registries such as npm, PyPI, Maven Central, and CRAN demonstrate the importance of structured metadata and explicit dependency graphs. Package manifests (e.g., package.json, POM.xml) encode versioning, licensing, and dependency constraints in machine-readable form, enabling automated resolution and reproducible builds (pypi-pep458, ; npm-trusted, ; hf-modelcards, ). SBOM standards require explicit declaration of components and dependency relationships for provenance and traceability (ntia-sbom-minimum, ; ntia-sbom-framing, ); emerging AI/ML BOMs extend the same idea to models, datasets, and configurations (xia2024trustsoftwaresupplychains, ; bennet2025SPDXAIBOM, ). Hugging Face relies primarily on model cards with limited dependency information (JiangPTMReuse, ). For example, while some cards reference required libraries, pretrained checkpoints, or paired datasets, these links are neither mandatory nor normalized into a dependency graph schema. This lack of standardized schemas leads to inconsistent naming and hampers automated reproducibility (Jiang2025Naming, ; JiangScored22, ; jiang2024peatmoss, ). The lesson for AgentHub is that agent metadata must go beyond identifiers to include standardized schemas that capture capabilities, input–output modalities, protocol requirements, and declared dependencies.
2.1.2. Provenance
@@
Trust and provenance mechanisms are central to registry design. Maven Central requires every artifact to be signed with a PGP key; npm supports registry signatures and “trusted publishing” using OIDC (npm-trusted, ; npm-signatures, ); and PyPI is adopting The Update Framework (TUF) for signed metadata (pypi-pep458, ). For agents, provenance is especially critical because dynamic tool bindings and environment access amplify the risks of impersonation or poisoning. Accordingly, AgentHub should require signed metadata and reproducibility attestations for all entries.
2.1.3. Governance & Submission Policies
@@
Governance models highlight trade-offs between openness and safety. Npm and PyPI accept broad participation with light pre-checks, while CRAN and app stores such as Apple’s App Store impose strict manual review  (apple-appstore-guidelines, ; npm-2fa, ). Ecosystems also implement revocation: app stores can remotely disable malicious apps, and PyPI can yank faulty releases (pypi-pep458, ). Notably, the leftpad incident in npm illustrated the ecosystem-wide disruption that can follow from a poorly governed removal (npm-leftpad, ; theregister-leftpad, ). Governance requires care and is non-obvious with autonomous agents.
2.1.4. Discovery & Quality Signals
@@
Registries, model hubs, and extension marketplaces provide user ratings, download counts, verification badges, and metadata-rich model cards (hf-modelcards, ; hf-popularity, ). These signals allow users to identify reputable contributions at scale. AgentHub should adopt reputation systems such as usage statistics, audits, or benchmark results to complement technical metadata.
@@ 3. Research Agenda
We distinguish challenges shared by all registries (section 3.1) from those unique to agents (section 3.2). We derived the AgentHub agenda as a scoped synthesis of two evidence streams: mature registry ecosystems for software and models, and emerging infrastructure for LLM-based agents. First, we reviewed the design and operational practices of widely used package and artifact registries (e.g., namespaces, versioning, provenance, governance, and security controls) alongside the software-engineering literature that analyzes their ecosystem dynamics and failure modes. In parallel, we examined recent agent directory and protocol efforts (e.g., agent cards and tool descriptors, agent-to-agent interaction protocols, and registry-like services) and extracted recurring friction points that arise when reuse becomes automated and cross-protocol. We then iteratively grouped these observations into a small set of agent challenges and mapped each challenge to the registry mechanisms required to mitigate it. The result is the set of six requirements in section 3.2, organized in Figure 2 to make the traceability from requirements to challenges and objectives explicit. This process is intended to be reproducible and extensible, so that as agent ecosystems mature, additional challenges can be incorporated and re-mapped without changing the core principle that registry mechanisms must be enforceable and machine-actionable to support autonomous reuse.
@@ 3.2. Agent-Specific Considerations
3.2.1. Discovery and Workflow Integration
Challenge: Keyword Search & Popularity Bias. Agents that recommend or install one another can create feedback loops that instill mediocre or unsafe entries. Experience from software registries highlights how naive signals mislead: Zerouali et al. show that different popularity metrics in npm yield inconsistent results (Zerouali2019, ), underlining how reliance on downloads or stars can mislead users; and Jones et al. observed that model popularity on HuggingFace correlates strongly with documentation quality (jones2024knowhuggingfacesystematic, ), suggesting that discovery in AgentHub would similarly benefit from incentives for high-quality documentation. Agents benefit greatly if those signals are captured as structured metadata and verifiable evidence. Without workflow integration, agents may fall back to ad-hoc search, undermining adoption and reproducibility.
Solution: AgentHub will improve current practice by combining keyword search with structured metadata and evidence linked to benchmarks, incorporating lifecycle states to avoid unsafe or outdated agents. Evaluation should identify which signals such as metadata and documentation quality ensure reliable discovery. Discovery should match user goals to agents by capabilities and evidence, ensuring that popularity does not overshadow quality. This resembles work on recommendation systems for software libraries like Code Librarian, which uses contextual signals to suggest packages (tao2023codelibrarian, ). Research benchmarking discovery is necessary to measure precision, recall, and resilience under ecosystem changes.
3.2.2. Trust and Security
Challenge: Privacy & Vulnerability. In software registries, account hijacks and malicious updates already erode trust (Zimmermann2019, ). LLM-based agents add privacy-specific attack surfaces: misconfiguration or weak governance can leak sensitive data or intellectual property (wang-etal-2025-unveiling-privacy, ; zharmagambetov2025agentdamprivacyleakageevaluation, ); training-data exposure can reconstruct confidential content (carlini2021extractingtrainingdatalarge, ); and real incidents have leaked corporate data (samsung_chatgpt_forbes_2023, ). Multi-agent prompt injection can propagate malicious instructions and compromise collective decision-making (lee2024promptinfectionllmtollmprompt, ; shi2025promptinjectionattacktool, ). Privilege escalation, guardrail bypasses, weak output validation, and insufficient access control enable unauthorized tool use and code execution (narajala2025securinggenaimultiagentsystems, ). The attack frontier remains underexplored.
Solution: AgentHub will deploy defenses such as signed manifests, provenance attestations (pypi-pep458, ; npm-trusted, ), verified namespaces via ANS (ans, ), and protections against typosquatting and package-confusion attacks (Zimmermann2019, ; jiang2025confuguardusingmetadatadetect, ). AI-specific risks include deserialization exploits and prompt-injection. It is crucial to test whether defenses from human-mediated registries hold for autonomous agent systems. New threats call for zero-trust privilege separation, runtime checks for I/O behavior, and privacy-preserving audit pipelines to verify evidence without leaks (he2025redteamingllmmultiagentsystems, ; narajala2025securinggenaimultiagentsystems, ; samsung_chatgpt_forbes_2023, ; wang-etal-2025-unveiling-privacy, ).
3.2.3. Openness & Governance
Challenge: Workforce & Concentration of Expertise. Many projects are maintained by a single person, creating bus-factor risk. Zimmermann et al. emphasize that the issue is less maintainer shortage than concentrated control (Zimmermann2019, ). Governance models differ: npm allows instant publication, while curated ecosystems like CRAN impose stricter checks.
Solution: Future work should investigate governance structures that combine openness with verifiability. Possible directions include decentralized namespace assignment rooted in ANS, community-driven policy boards for resolving disputes, and auditable logs of publication and revocation decisions. Comparative studies of centralized versus federated governance models can illuminate trade-offs in consistency, adoption, and resilience. Open questions include how to embed transparency in decision-making without sacrificing agility, and how to make governance mechanisms both fair across domains and enforceable at internet scale.
3.2.4. Ecosystem Interoperability
Challenge: One Schema, Many Dialects. Descriptors in MCP (tools), A2A (roles/behaviors), and ACP (message schemas/policies) use different primitives and evolve at different speeds. Without a standardized, signed manifest, semantics may be lost in translation, caches drift across mirrors, and naive popularity metrics dominate cross-protocol rankings while ignoring evidence quality and freshness. Stable identifiers for referenced artifacts are also missing, making cross-registry links fragile.
Solution: We propose a compact capability ontology and a canonical manifest with required fields (capabilities, I/O modalities, protocol bindings, permissions, SBOM-style dependencies) and optional per-ecosystem extensions. Declarative adapters can then map descriptors to this core and back, validated via round-trip conformance tests. Stable cross-registry identifiers (e.g., Purl-style) should be introduced and tested end-to-end with npm, PyPI, and model hubs. Early benchmarks could measure cross-protocol discovery in terms of precision/recall for intent queries, preservation of evidence link, freshness under churn, and ranking fairness as ecosystems scale.
@@ 3.2.5. Lifecycle Transparency
Challenge: Dependency Concentration & Abandonment. The leftpad incident showed how removing even trivial packages disrupted thousands of projects. Zimmermann et al. found that a small number of npm maintainer accounts control most packages (Zimmermann2019, ), while Zerouali et al. note that popularity often masks inactivity (Zerouali2019, ). Agents face parallel risks: if many depend on a single base agent or are generated by one entity, failure or compromise could cascade broadly. Abandonment may occur not only when humans leave but also when autonomous agents stop updating, leaving stale yet discoverable entries in circulation.